Calculate the percentage of water of crystallization in Washing Soda 

Grade:10

1 Answers

Sujit Kumar
5 years ago
Washing Soda (Hydrous)  = Na2CO3.10H2O
M.W of Na2CO3.10H2O = 2(23) + (12) + 3(16) + 10(18) =46 + 12 + 48 + 180 = 286
M.W of H2O in Na2CO3.10H2O = 10(18) = 180
 
\frac{w}{W}\%=\frac{180}{286}\times 100
\rightarrow \frac{w}{W}\%=0.62937\times 100
\rightarrow \frac{w}{W}\%=62.937\%
 
Ans: 62.9%